The change in internal energy for the combustion of 1.0 mol of octane at a pressure of 1.0 atm is 5084.5kJ . If the change in enthalpy is 5074.3kJ , how much work is done during the combustion?
Answer #1

Enthalpy is defined as H = U + pV, so
dH = dU + pdV + Vdp.

Assuming constant pressure of 1.0 atm, the final term vanishes.

Work is defined as dW = pdV, so
dH = dU + dW
dW = dH - dU

Thus,
dW = 5074.3 - 5084.5
dW = -10.2 KJ